Fermi National Laboratory
 

Operation Summary for September 19, 2011 - September 21, 2011

Monday September 19

The day shift began with Operations monitoring the Recycler stash (278.6E10), with the Antiproton Source (Pbar) stacking (20.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for a target change, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next Tevatron (TEV) store.

During the day shift, operators put the TEV through a dry squeeze at 8:25 AM. Operations established beam to the Meson FTBF experiment T-922 at 9:07 AM. Operators took the TEV separators out of conditioning mode at 9:23 AM. Cryo System personnel reported at 10:16 AM that the TEV sector F3 wet engine needed a complete overhaul, which would take about eight hours. Operations loaded the TEV at 10:37 AM with a 36x0 store for a study; the study ended at 10:46 AM due to a F3 lead failure that tripped off the TEV ramp. Operators put the TEV abort kickers in conditioning mode at 10:59 AM. A Pelletron person put the Pelletron through 30 minutes of conditioning at 11:26 AM. Fluids personnel reported at 1:21 PM that they were bringing column #2 online at CUB.

During the evening shift, operators prepared CDF for a two-hour access that began at 4:27 PM. Operations disabled beam to T-922 at 6:03 PM. Operators prepared D0 for a 20-minute access at 6:11 PM. Cryo System personnel reported at 6:50 PM that they had completed their F3 wet engine repairs. At 7:54 PM, the Cryo Coordinator reported bad vacuum on the F3 U-Tube. They had the U-Tube replaced by 9:26 PM. At 10:23 PM, the TEV LCW cut itself off from CUB due to a LCW resistivity problem. Operators called in Fluids personnel; they had the TEV and CUB operating together by 11:17 PM.

Tuesday September 20

The midnight shift began with the TEV turn on and recovery in progress, with Operations monitoring the stash (500.7E10), with Pbar stacking (124.9E10), and with MiniBooNE taking beam.

During the midnight shift, Operations put the TEV through a dry squeeze at 1:03 AM and through a wet squeeze at 2:27 AM. Operations put the TEV into HEP shot setup at 2:52 AM and established store 9139 at 3:50 AM with an average initial luminosity (AIL) of 355.93E30. Operations established beam to T-922 at 6:31 AM.

During the day shift, a MI person began an 8 GeV coalescing study at 9:35 AM. Operations established beam to T-922 at 10:02 AM. At 2:08 AM, a Linac quadrupole magnet power supply (QPS413) tripped off and would not reset; Linac technicians investigated and reported that they were beginning to replace the power supply. The MI person completed his coalescing study at 2:53 PM. Linac techs completed their work and Linac resumed sending beam at 3:23 PM.

During the evening shift, operators terminated store 9139 at 5:50 PM and put the TEV into shot setup. Operations disabled beam to T-922; this was the end of the experiment’s present run at Fermilab. At 6:55 PM, the TEV quenched ; operators investigated and found the TEV sector B1 vertical separator had sparked. Operators called in Separator personnel. At 7:47 PM, the MI-30 quadrupole power bus beam permit tripped off on a ground fault indication; it reset. Operators prepared the TEV for access at 8:49 PM to begin preparations for a Tevatron Electron Lens (TEL2) hollow beam study and to investigate problems with the TEV beam monitors. Operators reported short downtimes during the shift due to a Pbar AP-50 water skid trip, Linac quadrupole magnet power supply trips (QPS108 and QPS114), and to a Booster RF vacuum pump (B:IPRF17) trip.

Wednesday September 21

The midnight shift began with the TEV recovering from a quench, with Operations monitoring the stash (273.4E10) and the Pbar stack (11.9E10).

During the midnight shift, operators reported at 1:29 AM that the TEV quenched on its first ramp; operators investigated and believed the problem was due to lead flow failure. Operators reported that the TEV had recovered and at 4:10 AM they had put the TEV through a dry squeeze. Operators put the TEV through a wet squeeze and checked its orbits at 5:44 AM. Operations put the TEV into shot setup at 6:13 AM. At 6:45 AM, operators reported that a TEV RF station (TRF7) tripped off during the final proton load; operators called in a RF person.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct a study and for MI and Instrumentation personnel to load new software as a test for the MI BPM system.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/21/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory